### v1

# HPLT 3.0: Very Large-Scale Multilingual Resources for LLM and MT Mono- and Bi-lingual Data, Multilingual Evaluation, and Pre-Trained Models

Abstract
We present an ongoing initiative to provide open, very large, high-quality, and richly annotated textual datasets for almost 200 languages. At 30 trillion tokens, this is likely the largest generally available multilingual collection of LLM pre-training data. These datasets are derived from web crawls from different sources and accompanied with a complete, open-source pipeline for document selection from web archives, text extraction from HTML, language identification for noisy texts, exact and near-deduplication, annotation with, among others, register labels, text quality estimates, and personally identifiable information; and final selection and filtering. We report on data quality probes through contrastive and analytical statistics, through manual inspection of samples for 24 languages, and through end-to-end evaluation of various language model architectures trained on this data. For multilingual LLM evaluation, we provide a comprehensive collection of benchmarks for nine European languages, with special emphasis on natively created tasks, mechanisms to mitigate prompt sensitivity, and refined normalization and aggregation of scores. Additionally, we train and evaluate a family of 57 monolingual encoder–decoder models, as well as a handful of monolingual GPT-like reference models. Besides the monolingual data and models, we also present a very large collection of parallel texts automatically mined from this data, together with a novel parallel corpus synthesized via machine translation.
Keywords: large language models, multilinguality, pre-training data, evaluation, machine translation, mining parallel texts, synthetic data
[…]
HPLT 3.0: Very Large-Scale Multilingual Resources for LLM and MT
Mono- and Bi-lingual Data, Multilingual Evaluation, and Pre-Trained Models
Stephan Oepen, Nikolay Arefev, Mikko Aulamo, Marta Bañón, Maja Buljan,
Laurie Burchell, Lucas Charpentier, Pinzhen Chen, Mariya Fedorova, Ona de Gibert,
Barry Haddow, Jan Hajič, Jindrič Helcl, Andrey Kutuzov, Zihao Li, Risto Luukkonen,
Bhavitvya Malik, Vladislav Mikhailov, Amanda Myntti, Dayyán O’Brien, Lucie Poláková,
Sampo Pyysalo, Gema Ramírez Sánchez, Janine Siewert, Pavel Stepachev,
Jörg Tiedemann, Teemu Vahtola, Fedor Vitiugin, Tea Vojtěchová, Jaume Zaragoza
University of Oslo, Department of Informatics
University of Helsinki, Department of Digital Humanities
Prompsit Language Engineering
The Common Crawl Foundation
Edinburgh University, School of Informatics
Charles University, Prague, Institute of Formal and Applied Linguistics
University of Turku, Department of Computing
oe@ifi.uio.no
Abstract content
1.  Introduction
Massive text collections for pre-training are the “crude oil” of the LLM era. The process of “refining” high-quality datasets from web data at scale presupposes computational infrastructure and technological muscle that oftentimes is characteristic of corporate involvement, as evidenced for example by some notable generally available pre-training datasets: C4 (the Colossal Clean Crawled Corpus, by Google and Allen AI; Raffel et al., 2020), FineWeb 1 & 2 (by Hugging Face; Penedo et al., 2024, 2025), MADLAD-400 (by Google; Kudugunta et al., 2023), or Nemotron-CC (by Nvidia; Su et al., 2025). With notable exceptions, this line of work tends to capitalize on the English language.
We build on the open-source pipeline of the European academic consortium HPLT (High-Performance Language Technologies; de Gibert et al., 2024; Burchell et al., 2025), a project that was funded under the Horizon Europe programme in 2022–2025. We revise and improve various of the sub-components, and apply the updated pipeline to a greatly enlarged collection of raw web data. This results in a massive pre-training dataset of high-quality texts in close to 200 distinct language–script combinations. We dub this novel resource HPLT 3.0. The data comprises some 30 trillion sub-word tokens in total, of which close to half represent languages other than English.
Following the examples of FineWeb and HPLT 2.0, we make this resource publicly available under permissive licenses. [^1] https://hplt-project.org/datasets/v3.0 We further share the refined open-source processing pipeline, a novel multilingual evaluation framework, as well as various collections of language models pre-trained on HPLT 3.0 data. Finally, while our focus in this report is on the monolingual data and models, we also briefly report on ongoing work to derive novel bilingual datasets for 28 language pairs, provide associated machine translation models, and synthesize additional pre-training data for underrepresented languages by machine translation of very high-quality English documents. In our view, it is the totality of generally available and very large-scale resources and the documentation of the underlying processes that bears promise of “democratizing” the current LLM and MT landscape. In-depth technical and experimental details will be provided through appendices upon publication.
2.  Raw Web Archive Data
There are few available collections of massive web archives. Our work builds on the same set of so-called “wide crawls” from the Internet Archive (IA) as the HPLT 2.0 release, but combines this with a broader and much larger set of snapshots from the Common Crawl (CC). Specifically, we start from some 3.3 petabytes of IA crawls from the period 2012–2020. We complement this data with 57 CC full snapshots from the period 2014–2025, making sure to include all available data since 2020, together with about half of the earlier CC snapshots. This amounts to almost five times more raw CC data than in HPLT 2.0, bringing our total volume of web archives to about 7.2 petabytes. HPLT 3.0 to date does not include a few experimental data sources, an IA “survey crawl” and a minor sample from the ArchiveBot repository (270 terabytes in HPLT). We plan on reporting on an enlarged dataset in early 2026, augmented with another 3 petabytes from ArchiveBot.
3.  Monolingual Data Preparation
Figure 1: Schematic overview of data preparation.
Extraction of high-quality and richly annotated text from raw web archives proceeds through a sequence of refinement and filtering steps. Figure 1 shows the main components of our data preparation, which is a refined and extended version of the open-source HPLT pipeline by Burchell et al. (2025). The following paragraphs provide a high-level overview of the refinements we make to the original HPLT 2.0 pipeline.
Text Extraction
Like FineWeb and HPLT 2.0, we stick to the Trafilatura text extraction framework Barbaresi (2021), but we conduct a comprehensive data-driven optimization of its many hyper-parameters, emphasizing extraction quality over speed and text recall.
Language Identification
We predict the language of the text during preprocessing using OpenLID-v2, an updated version of the model described in Burchell et al. (2023). Compared to the prior version of OpenLID, we revise the inventory of language labels to align with the Flores+ evaluation set Burchell et al. (2024), and include additional training data for Dari, South Azerbaijani, Asturian, and Paraguayan Guaraní. We also revise input preprocessing for language identification to a simpler, language-agnostic pipeline: We normalize whitespace, apply lowercasing, and remove non-word characters and digits. This change increases the robustness of language identification to the non-standard orthography commonly found in web documents.
Deduplication
Random repetitions caused by overlapping crawls can negatively impact LLM training Lee et al. (2022). While HPLT 2.0 followed the original FineWeb design and limited itself to per-collection or per-crawl deduplication, we implement MinHash-based global near-deduplication for all languages in HPLT 3.0 except English, Russian, and Chinese for which we stick to per-crawl deduplication for computational efficiency.
Annotation
The Web Docs Scorer (WDS) approach [^2] https://github.com/pablop16n/web-docs-scorer provides an integrated consolidation of a rich tradition of heuristic document filtering, e.g. based on length statistics, presence of “oddity” signals, and document- vs. segment-level language distributions. We adapt and refine this tool and make WDS levels a central annotation in the HPLT 3.0 dataset (see below). Furthermore, we update and re-train the Turku web register classifier Myntti et al. (2024) and apply register annotations for 104 of the languages in HPLT 3.0.
HPLT 3.0 FineWeb 1.4.0, 2.1.0 HPLT 2.0 MADLAD-400 1.0
Language D T | | % D T | | % D T | | % D T | | %
English 18B 16T 901 55 24B 17T 695 78 4.4B 3.9T 892 35 1.5B 1.7T 1099 38
Multilingual 11B 13T 1187 45 5.0B 4.9T 976 22 6.1B 7.2T 1178 65 2.1B 2.7T 1266 62
Basque 3.2M 3.2B 991 0.02 1.6M 1.5B 951 0.03 2.0M 2.0B 1030 0.03 1.6M 1.5B 1258 0.05
Catalan 2.6M 22B 853 0.17 1.7M 12B 715 0.25 1.9M 18B 976 0.25 0.9M 10B 1084 0.38
Czech 107M 126B 1171 0.93 66M 67B 1015 1.37 75M 95B 1266 1.32 38M 50B 1329 1.88
Finnish 49M 73B 1491 0.55 36M 48B 1324 0.99 34M 53B 1538 0.74 20M 34B 1673 1.27
French 603M 584B 968 4.32 360M 292B 811 5.95 401M 379B 943 5.24 216M 269B 12403 9.95
Galician 4.0M 3.1B 772 0.02 2.5M 1.8B 695 0.04 3.0M 2.7B 906 0.04 0.2M 1.3B 1004 0.05
Norwegian 37M 52B 1388 0.39 40M 53B 1318 1.09 28M 42B 1477 0.58 14M 22B 1508 0.83
Spanish 725M 658B 908 4.86 441M 329B 746 6.71 503M 471B 936 6.51 250M 254B 1015 9.43
Ukrainian 80M 81B 1014 0.60 53M 49B 938 1.02 47M 60B 1280 0.84 24M 31B 1268 1.17
Table 1: Key statistics for the HPLT 3.0 data, contrasted with FineWeb, HPLT 2.0, and MADLAD-400. Columns show document and token counts, as well as average document length (“| |”) and the token share of each subset of the total (“%”); individual per-language percentages are of the non-English parts only.44footnotemark: 4
Packaging
The WDS scores described above provide a concise and language-agnostic perspective on document properties that we expect to correlate with their prospective utility in LLM development. For flexible experimentation with WDS-based data sampling from HPLT 3.0, we bin documents within each language by WDS levels – $\{$5, 6, 7, 8, 9 10}\} – and globally sort each bin. § 7.1 below presents experimental evidence that training on higher WDS levels can enhance LLM performance. All HPLT 3.0 data – documents and metadata – is packaged in Zstandard-compressed JSONlines form, where larger WDS bin are broken up into multiple shards. The total release comprises some 50 terabytes in about 3000 files, which are distributed via public HTTP download.
4.  Overall Dataset Statistics
To put the HPLT 3.0 monolingual dataset into perspective, Table 4 presents document and token counts [^5] For the purpose of comparable statistics across languages and different datasets, all token counts are computed using the Gemma-3 tokenizer Team (2025), a SentencePiece-based tokenizer with a vocabulary of 256K sub-words, providing good coverage for all target languages. for the English and multilingual (‘non-English’) partitions of the data, as well as counts for a small sample of individual languages. For ease of comparison, these statistics are accompanied with average document lengths and per-language proportions, and contrasted with corresponding figures for three other publicly available multilingual datasets mentioned in § 1. 55footnotetext: All dataset references in the remainder of this paper are to the specific versions indicated in the table headers.
As is evident from these numbers, HPLT 3.0 is the by far largest publicly available such dataset, and its multilingual breadth compares favorably to other widely used resources. In Gemma-3 tokens, the multilingual HPLT 3.0 partition is about 2–3 times larger than FineWeb and HPLT 2.0, respectively, and five times larger than the older MADLAD-400 dataset. In terms of average document length, HPLT 3.0 patterns with HPLT 2.0, markedly ahead of FineWeb and well behind MADLAD-400. A small selection of European languages in Table 4 shows languages ranging between a “mere” billion of available tokens to others with hundreds of billions.
5.  In-Depth Analytics
Like Burchell et al. (2025), we calculate descriptive statistics for HPLT 3.0 using the HPLTAnalytics tool [^6] https://github.com/hplt-project/data-analytics-tool and compare HPLT 3.0 to their dataset.
Descriptive Statistics
We notice a substantial difference in unique segments, 73% in HPLT 3.0 vs. 52% in HPLT 2.0, on average. This likely reflects global rather than per-crawl deduplication (see above) that removes near-identical documents, thus increasing text diversity. The difference is in general higher for small-to-medium languages and lower in larger datasets, which generally tend to have a smaller proportion of unique segments. However, the ratios of large documents (above 25 segments) and short segments (less than 3 tokens) show no significant differences: 21% vs. 20% for large documents and 13.1% vs. 11.5% for short segments (on average), for HPLT 3.0 and 2.0, respectively. Similarly, proportions of individual segments in the document language are comparable, at 71.5% vs. 71.9%, on average. Lower values in this metric are usually associated with low resourced languages or languages easily confused with similar, more-resourced ones.
Internet Domains
Compared to HPLT 2.0 and because of the global deduplication process, HPLT 3.0 exhibits a wider variety of domains. For example, we notice that while HPLT 2.0 has more than 20 datasets with at least 25% of their documents coming from Wikipedia pages, this holds for only seven datasets in HPLT 3.0 (Wikipedia documents in HPLT 2.0 are likely to be repeated across collections, thus overrepresented in the dataset). However, and similar to HPLT 2.0, biblical webpages are a large part of the content of smaller language datasets (especially African languages), with up to 25 datasets in HPLT 3.0 with more than half of their documents originating from this kind of domains.
Regarding Top Level Domains (TLDs), we find that the geographic TLD distribution in HPLT 3.0 is similar to HPLT 2.0, usually with the most frequent geographic TLD corresponding to the country where the language of the dataset is spoken. This ratio tends to be higher in mid-sized European languages. Similarly, we find languages containing geographic TLDs from various countries where its language is spoken (suggesting the proportions of different variants of the language in the dataset) and some datasets showing a variety of geographic TLDs from closely related countries or territories, indicating possible deficiencies with language identification.
N-grams
HPLTAnalytics calculates the five most common $n$-grams for orders 1 to 5 in each dataset, after discarding $n$-grams that start or end with a stopword. Regarding frequent $n$-grams, the most pronounced difference when comparing to HPLT 2.0 is the lack of Wikipedia-related $n$-grams (see above). We also find indicators of low-quality text in some of the datasets that are more prominent in HPLT 3.0 than in 2.0, such as adult (notably in European languages) or betting content.
Register Labels
After inspecting the distribution of web registers (see above) among the different languages of HPLT 3.0 for which they are available, we find that the most common register tends to be Narrative, especially the News Report subregister. However, for larger European languages, the Narrative register tends to be tied or surpassed by Information Persuasion, noticeably the Description with Intent to Sell subregister. In HPLT 2.0, the same two registers appear the most frequent, but in this case the most frequent subregister labels are Narrative Blog and Encyclopedia Article. This contrast in blogs and encyclopedic articles documents is likely also attributable to global deduplication.
6.  Manual Inspection
For 23 languages, native or fluent speakers have manually inspected randomly sampled documents and marked those that contain pornographic content, text with artifacts (e.g. navigational elements, headings or list items without proper delimitation, truncated text, or snippet markers), unnatural text (e.g. word lists for search engine optimization, high proportions of “boilerplate”, or very obvious machine translation), or incorrectly identified language. Reflecting availability of annotators, for each language between 50 and 1000 documents were inspected, on average 348 documents per language. Table 2 reports confidence intervals for percentages of problematic texts of different types.
Language Porn Artifacts Unnatural LID
Asturian 0-1 2-9 2-8 19-31
Bosnian 0-0 52-62 3-8 66-75
Catalan 0-1 0-3 0-1 0-2
Chinese 0-3 0-5 2-8 0-2
Croatian 0-2 42-52 07-14 06-12
Czech 0-1 4-7 24-30 0-1
English 0-2 29-43 0-5 0-1
Finnish 2-4 15-20 07-11 0-0
French 0-4 13-23 10-21 0-1
Galician 0-1 04-11 0-5 0-3
German 0-1 0-3 05-13 0-1
Hindi 0-1 20-33 1-6 0-4
Iranian Persian 0-1 35-45 08-15 1-4
Italian 0-4 04-11 05-14 0-2
Japanese 0-2 47-61 13-23 0-1
Modern Greek 0-1 31-37 2-5 0-1
Norwegian Bokmål 2-7 12-20 2-7 0-3
Portuguese 0-2 30-44 04-11 0-3
Russian 0-1 4-9 2-5 0-1
Serbian 0-6 48-75 06-27 0-6
Slovak 0-0 11-19 4-9 0-2
Spanish 0-1 15-23 05-10 0-1
Yoruba 0-3 07-16 0-5 04-11
Table 2: Manual inspection of the HPLT 3.0 dataset samples: 95% confidence intervals for the percentages of texts containing porn, artifacts, unnatural texts, and texts in a different language (LID errors).
We observe that the proportion of porn in the dataset is below 2% for most languages. The proportion of the documents where the language is mis-identified is also quite low, with the notable exceptions of the Bosnian dataset, which consists mostly of texts in Serbian, and the Asturian dataset, which often contains Spanish. The proportions of unnatural texts and texts containing artifacts widely vary across languages. This is related in part to the fuzzy definition of these properties in the guidelines and subjectivity in judgments. For example, among two annotators of Czech, one annotated 34% of text as unnatural while another only 1%. The notion of text naturalness is subjective by definition.
7.  Multilingual LLM Evaluation
We develop HPLT 3.0-E, a framework for automated large-scale multilingual evaluation designed to systematically compare and refine data preparation choices across nine selected languages shown in footnote 4. These languages are chosen to ensure both availability of native speakers in our development team and a minimum level of diversity in terms of language resources, families, and scripts. Similar to Penedo et al. (2024, 2025), we pretrain separate language models per language using an otherwise fixed pretraining setup, and evaluate them at regular checkpoint intervals (every 1B tokens) in a 0-shot regime, carefully selecting tasks that meet the pretraining evaluation signal criteria below.
7.1.  HPLT 3.0-E: Framework
HPLT 3.0-E includes 127 language understanding and generation tasks, each supporting 3–7 human-written prompts. We aim to cover different task categories in all languages: entailment, commonsense reasoning, language-specific & world knowledge, paraphrasing, reading comprehension, sentiment analysis, toxicity detection, and truthfulness. HPLT 3.0-E integrates with LM Evaluation Harness Gao et al. (2024), for experimental flexibility and replicability.
Pretraining
We pretrained decoder-only models of the size 2.15B on 100B tokens sampled from FineWeb, HPLT 2.0, MADLAD-400, and HPLT 3.0.777For lower-resource languages with less than 100B tokens of available data, datasets are uniformly upsampled (repeated) following Muennighoff et al. (2023a). All models employ the Gemma-3 tokenizer and follow the Llama architecture Touvron et al. (2023) with 24 layers, 32 attention heads, and a sequence length of 2048. Pretraining is run using the Megatron-LM Shoeybi et al. (2019) framework on the LUMI supercomputer, utilizing 16 nodes with AMD MI250x GPUs for a total of approximately 3,000 GPU hours per model. The estimated carbon footprint per model is 59 kg CO2.
Benchmarks
We combine open-source human-curated multi-task benchmarks: IberoBench (Catalan, Spanish, Basque, Galician; Baucells et al., 2025), FrenchBench (French; Faysse et al., 2024), NorEval (Norwegian Bokmål and Nynorsk; Mikhailov et al., 2025), BenCzechMark (Czech; Fajcik et al., 2025), and Finbench v2 built on Finbench (Finnish; Luukkonen et al., 2023). In addition, we create a benchmark for Ukrainian (UkrainianBench), which comprises Global MMLU Singh et al. (2025), INCLUDE Romanou et al. (2025), UA-SQuAD Ivanyuk-Skulskiy et al. (2021), ZNO Romanyshyn et al. (2024), Belebele Bandarkar et al. (2024), TextDetox Dementieva et al. (2024), and WMT24++ Deutsch et al. (2025b).
Prompt Collection
HPLT 3.0-E enables evaluation across 500+ prompts that have diverse structure and answer formatting to mitigate prompt sensitivity, a model limitation where variations in prompt formulation can affect downstream performance (e.g. Pezeshkpour and Hruschka, 2024; Sclar et al., 2024). We adapt single-prompt benchmarks (IberoBench, FrenchBench, and UkrainianBench) to the multi-prompt design through (i) manual translation of English prompts from PromptSource Bach et al. (2022) and (ii) prompt creation by native speakers.
Task Selection
We use standard task-specific metrics and report the maximum score across the prompts as the main aggregation method. We extend the FineTasks evaluation design Penedo et al. (2025) and select tasks that provide pretraining evaluation signal based on the following key criteria: monotonicity – the model performance should improve as pretraining progresses; stable pretraining – the relative variability of model performance across pretraining intervals should be low; non-randomness – the final model performance should exceed a random guessing baseline and be satisfactory; ranking consistency – the relative ranking of models across pretraining intervals should remain stable; low noise – the final model performance should be high relative to task variability across prompts; low prompt sensitivity – the median absolute deviation across prompts should be low to ensure prompt-invariant dataset comparison; prompt lottery – the frequency with which the best-performing prompt changes as pretraining progresses should be low.
Performance Aggregation
Following Fourrier et al. (2024); Penedo et al. (2025), we compute a language score as the average of min–max-normalized performance scores across selected tasks. In particular, we first rescale all scores between the random baseline and the maximum possible score. We then average the rescaled scores within each task category and take the average of per-category scores to compute the language score. To produce a multilingual score, we utilize several approaches: average language score – we compute the average of all language scores; average multilingual rank – we rank the final models’ language scores across all datasets and average their ranks; Borda’s count – using Vote’n’Rank Rofin et al. (2023), we compute the final models’ Borda rankings within each language and aggregate the per-language rankings to produce the overall ranking. Borda’s count serves as an alternative to average-based aggregation, allowing for aggregating heterogeneous metrics by leveraging rank-based differences among models.
7.2.  Results
In line with Penedo et al. (2025), we find that tasks for lesser-resourced languages, notably Basque and Galician, are unsuitable for pretraining evaluation due to potential difficulty and lack of monotonic performance progression during pretraining. We thus report our key findings on a final suite of 27 selected tasks across seven remaining languages.
Dataset Comparison
Figure 2 presents the results of comparing the models pretrained on different datasets. All models show a monotonic performance improvement on our selected tasks as pretraining progresses. Models pretrained on MADLAD-400 achieve the highest multilingual score, followed by HPLT 3.0, while HPLT 2.0 and FineWeb perform on par.
These results are consistent with rank-based aggregation. Models are ranked as (1) MADLAD-400; (2) HPLT 3.0; and (3) HPLT 2.0 and FineWeb; by average multilingual ranks, HPLT 2.0 slightly outperforms FineWeb, whereas Borda’s counts show the inverse ordering. Overall, our findings indicate that refined data preparation in HPLT 3.0 has improved average dataset quality, which translates into competitive performance gains for model pretraining. The performance of models trained on the older and smaller MADLAD-400 data calls for follow-up studies.
Figure 2: Comparison of models pretrained on FineWeb, HPLT 2.0, 3.0, and MADLAD-400.
Sampling by Quality Estimates
As another application of HPLT 3.0-E, Figure 3 shows the performance of three Spanish models trained on different subsets of HPLT 3.0 training data selected according to WDS levels; see § 3 above. In these experiments, five Spanish tasks were selected as informative for all models. Here, “random” sampling represents the default approach, drawing uniformly on the full dataset, while “top” and “bottom” take advantage of the sorting by WDS levels and sequentially draw 100B training tokens from either end of the dataset. Low WDS levels clearly lead to inferior model performance, while sampling from only the “top” does not cearly improve over the full dataset, possibly owing to overly limited diversity. We observe similar patterns in French, the other language supported by HPLT 3.0-E where there are hundreds of billions of available training data.
8.  Monolingual Encoder–Decoders
In this section, we apply the HPLT 3.0 dataset to train and evaluate 57 language-specific monolingual encoder–decoder language models, following the T5-base architecture Raffel et al. (2020).
Figure 3: Comparison of different WDS-based sampling strategies from the Spanish HPLT 3.0 data.
Motivation & Approach
Despite the popularity of decoder-only LLMs in recent years, encoder–decoders are still widely used in real-world applications, showing strength in both generative and discriminative tasks Zhang et al. (2025).
Our T5-like models serve two primary purposes:
  1. 1.
    to evaluate HPLT 3.0 quality as training data across a large number of languages; and
  2. 2.
    to provide a family of comparable monolingual encoder–decoders trained on current data.
As regards the second purpose, we note the only available encoder–decoder with multilingual capabilities is mT5 pretrained on the mC4 corpus Xue et al. (2021), and its instruction-tuned derivatives, e.g. Muennighoff et al. (2023b); Üstün et al. (2024). Our collection of T5-like models is trained on more recent HPLT 3.0 datasets, which we believe to be substantially cleaner and more high-quality than mC4 (see our evaluation below). Thus, we hope that the NLP community will benefit from these novel models for the 57 languages that we currently support. In addition, language-specific models can be used to conduct comparative studies of LLM learning process (we also publish intermediate checkpoints). Their size ($\approx$275M parameters each) supports for easy and computationally cheap deployment. Our models follow the setup of Samuel et al. (2023).
Selection of Languages
Our choice of languages to train encoder–decoders on was motivated by their typological diversity. This means we aimed at covering as many different language families as possible. At the same time, we did not train models on extremely under-represented languages (less than $\approx$0.25M documents in our datasets).As a result, our set of 57 models covers the following language families: Indo-European, Sino-Tibetan, Japanese, Austronesian, Austro-Asiatic, Uralic, Altaic, Afro-Asiatic, Korean, Tai-Kadai, Dravidian, Kartvelian, Niger-Kongo, and Basque.
Language Named entity recognition (WikiAnn, F1) Linguistic competence (MultiBLIMP, accuracy)
size BERT HPLT T5 HPLT 3.0 mT5-base size T5 HPLT 3.0 mT5-base mT5-xxl
Catalan (cat_Latn) 10000 94.5 92.7 87.4 2284 95.6 91.6 93.0
Czech (ces_Latn) 10000 91.8 91.6 85.2 4256 95.9 88.8 93.4
English (eng_Latn) 10000 82.7 82.1 77.6 770 94.2 90.6 95.3
Basque (eus_Latn) 10000 92.9 92.0 82.8 273 97.4 94.9 96.0
Finnish (fin_Latn) 10000 91.6 90.3 01.8 2570 95.6 81.4 86.1
French (fra_Latn) 10000 90.0 88.9 83.3 2548 93.6 91.7 94.8
Galician (glg_Latn) 10000 94.1 93.4 89.2 753 96.0 90.7 95.4
Bokmål (nob_Latn) 10000 93.2 91.8 87.0 *3463 40.6 68.0 71.8
Nynorsk (nno_Latn) 1000 95.5 94.0 88.2 - - - -
Spanish (spa_Latn) 10000 90.8 90.7 84.0 2541 95.2 93.8 96.3
Ukrainian (ukr_Cyrl) 10000 92.8 92.5 82.1 2744 95.7 89.4 94.8
Average - 90.5 90.5 78.8 - 93.5 86.8 91.4
Table 3: Evaluation results of HPLT 3.0 monolingual encoder–decoders (11 languages of interest, Bokmål and Nynorsk are two varieties of Norwegian), along with the test set sizes for each language. Average results are computed over all the 57 languages we have trained models for (full results will be presented in the appendix upon acceptance). *Bokmål competence benchmarks are not part of MultiBLIMP.
Evaluation Results
We evaluate our T5-like models on two tasks: (i) named entity recognition (NER) using the WikiAnn benchmark Rahimi et al. (2019), (ii) linguistic competence using the MultiBLiMP benchmark Jumelet et al. (2025). Their performance was compared to the performance of the original mT5-base [^8] https://huggingface.co/google/mt5-base model and the language-specific BERT models [^9] https://huggingface.co/collections/HPLT/hplt-20-bert-models trained on the HPLT 2.0 datasets. For reference, we also demonstrate the performance of a much larger mT5-xxl [^10] https://huggingface.co/google/mt5-xxl model on the MultiBLiMP benchmark.
MultiBLIMP is a dataset of minimal pairs: each sample consists of a pair of sentences. The first sentence is grammatical (‘correct’). The second sentence is the same as the first one, but its syntactic head is modified in such a way, that the resulting sentence is ungrammatical (‘wrong’). We re-use MultiBLIMP’s official evaluation code, adding support for encoder–decoder models. Regardless of the particular model architecture, evaluation works as follows: probabilities of both ‘correct’ and ‘wrong’ sentences are produced using the language model under evaluation. The resulting accuracy is the share of samples where the probability of the ‘correct’ sentence is higher than that of the ‘wrong’ one. The difference between decoder-only and encoder–decoder models is that with the former, each token is only conditioned on the previous tokens and a sentence may be evaluated ‘as is’, while with the latter, each token is conditioned both on the encoder’s output and the previous tokens; thus, we replicate the T5 training procedure and mask the syntactic head in the encoder input, and the rest of tokens in the decoder input.
Since Norwegian is missing from MultiBLIMP, we used the grammatical error correction benchmark NoCola Jentoft and Samuel (2023) for Bokmål. Unlike MultiBLIMP, it may contain more than one ungrammatical sequence (derived from real Norwegian as a second language learners’ mistakes), and it may be not only words, but also punctuation marks, e.g. a single comma. One might expect that a model trained on web data would perform poorly on such a dataset, and it turned out to be true: Bokmål is the only language, for which both mT5 models achieve higher accuracy than our model.
Table 3 shows the selected evaluation results for the languages of interest as defined in § 7 plus English. It demonstrates that the HPLT 3.0 monolingual encoder–decoders offer a competitive alternative to the multilingual mT5 models. On average, our models achieve the same performance as HPLT BERTs on the NER task, while at the same time possessing all the advantages of encoder–decoder models in comparison to encoder-only architectures; they also outperform both mT5 models on the MultiBLIMP task.
We also conducted additional evaluations on the English MultiBLIMP: The original monolingual T5-base performed better (93.5% accuracy), than the multilingual one; instruction-tuned derivatives mt0-xl (90.5% accuracy) and aya-101 (86.6% accuracy) further proved the finding by Jumelet et al. (2025) that fine-tuning worsens a model’s BLIMP performance; the most recent English T5-base model t5gemma-b-b-ul2 showed surprisingly low result (66.5% accuracy); however, its training objective was more sophisticated Tay et al. (2023) than in the original T5. We leave further adjusting our inputs to it for future work.
9.  Mining for Bilingual Texts
After constructing and evaluating the monolingual corpus, a natural next step is to further leverage these resources to mine parallel data. Although the field is increasingly favoring LLMs over traditional encoder–decoder architectures, the use of parallel corpora in LLM pretraining has been shown to enhance multilingual capabilities Zhang et al. (2024); Qorib et al. (2025). Moreover, machine translation (MT) systems can be used to generate synthetic data from monolingual resources that is useful for LLM pretraining, as presented in § 10.
We are releasing parallel data at both sentence and document-level. Recent work in MT is moving towards document-level translation. This shift is reflected in the release of training datasets Merx et al. (2025); O’Brien et al. (2025), dedicated benchmarks Fernandes et al. (2023); Deutsch et al. (2025a); and the development of targeted training strategies Ramos et al. (2025). This is motivated by the ability of LLMs to leverage long context, which allows them to model discourse phenomena such as coreference resolution, anaphora or ellipsis, which are typically lost when restricting training to the sentence level.
We create English-centric parallel resources for 28 European languages, diverse in terms of linguistic families and resource availability. We use Bitextor [^11] https://github.com/bitextor/bitextor for parallel data extraction. Bitextor takes as input the cleaned monolingual documents. The pipeline performs sentence segmentation, translates sentence into synthetic English for document alignment, and then applies cleaning rules and deduplication. We follow de Gibert et al. (2024); Burchell et al. (2025) with minor modifications. The main change we apply is that we maintain document structure from the beginning of the extraction.
10.  MT for Synthetic Data Generation
Many studies have shown the value of synthetic data, for example, Doshi et al. (2024); Wang et al. (2025) and in this work we explore the use of machine translation as an effective short-cut to transfer knowledge from a resource-rich source language to a under-resourced target languages de Gibert et al. (2025).
Many open-source models are available and can easily be integrated in translation workflows Tiedemann et al. (2023); Team et al. (2022). For scalability, inference-efficient models are needed and, for that reason, we rely on compact encoder–decoder models instead of instruction-tuned LLMs. In particular, we select high-performing models available through the OPUS-MT Dashboard Tiedemann and de Gibert (2023) and sample around 28 billion tokens of English data from FineWeb-Edu Lozhkov et al. (2024) and 100 billion tokens from Nemotron-CC Su et al. (2025) (taking the high-quality subset) to be translated. We split documents into sentences using the LoomchildSegmenter [^12] https://pypi.org/project/loomchild-segment/. Short segments are merged into multi-sentence segments until they exceed a maximum threshold of 1,024 characters to improve context-sensitivity. We translate in shards of 500M tokens using Marian-NMT Junczys-Dowmunt et al. (2018) and beam size four. After translation, we merge translated sentences back into the original document structures, leading to a document-level and segment-level aligned parallel dataset. We plan to publish the synthetic data under a permissive license in the near future.
FineWeb-Edu translations are available in 36 languages and language variants, [^13] Language coverage (ISO-639-3 codes): bos, bul, cat, ces, dan, deu, ell, est, eus, fin, fra, gle, glg, hrv, hun, isl, ita, kat, lav, lit, mkd, mlt, nld, nno, nob, pol, por, ron, slk, slv, spa, sqi, srp, swe, tur, ukr. creating a corpus of around 1 trillion tokens. Translations from Nemotron-CC currently comprise 27 languages and 2.5 trillion tokens in total with additional languages in the loop to match the same language coverage as FineWeb-Edu translations.
We conduct initial evaluations of the translated Nemotron-CC data through training “smallish” decoder-only models for a comprehensive subset of languages. In these experiments, reference models with two billion parameters exclusively trained on synthetic data performed on par or even better than comparable models trained on cleaned “native” web data, testing with common multilingual benchmarks. These encouraging results indicate the potential value of the translated data, but also call for more comprehensive studies – along the lines of § 7, emphasizing language quality benchmarks, and preferably at larger scales – are needed to scrutinize these findings further.
11.  Conclusions & Outlook
Our work substantially refines an existing pipeline for very large-scale preparation of mono- and bi-lingual datasets and applies it to a massive collection of web archives. All data, models, and software involved in this initiative are publicly available under permissive terms of use. The HPLT 3.0 dataset constitutes the largest multilingual resource of its type. Contrastive experimentation suggests that its overall data quality enables LLM performance superior to comparable resources. This resource is supported by a novel multilingual evaluation framework and accompanied by multiple families of pre-trained language models for a broad range of languages, as well as by derived bilingual and synthetic text data. Through this work, we hope to contribute to increased community activity emphasizing generally available and transparent resources and processes for very large-scale LLM research, with special attention to multilinguality and linguistically and culturally adapted evaluation. Upon publication, we expect to invite community contributions to futher development, utlization, and refinement of these resources.
12.  Ethics Statement
Datasets crawled from the Web can contain all sorts of offensive or harmful content, even though we are making our best to clean the data. Datasets for lower-resourced languages can be biased in very different ways (e.g. religious content is often over-represented in African languages datasets).
13.  Limitations of this Work
Although we conducted a (limited) human evaluation of a selection of languages from the HPLT 3.0 datasets, the problem of robustly estimating the quality of web-crawled training data is far from being solved. We do not claim that the datasets described above are completely free from defects.
14.  Acknowledgments
We thank Étienne Simon (UiO) and Daryna Dementieva (TUM) for their contribution to our prompt collection for French and Ukrainian.
This project has received funding from the Horizon Europe research and innovation programme of the European Union under Grant agreement No. 101070350 and from UK Research and Innovation (UKRI) under the UK Horizon Europe funding guarantee, grant number 10052546. The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the European Union. The authors wish to thank CESNET, Czech Republic, CSC, Finland, and Sigma2, Norway, for computational resources and support.
15.  Bibliographical References
[…]
  • Bach et al. (2022) Stephen Bach, Victor Sanh, Zheng Xin Yong, Albert Webson, Colin Raffel, Nihal V. Nayak, Abheesht Sharma, Taewoon Kim, M Saiful Bari, Thibault Fevry, Zaid Alyafeai, Manan Dey, Andrea Santilli, Zhiqing Sun, Srulik Ben-david, Canwen Xu, Gunjan Chhablani, Han Wang, Jason Fries, Maged Al-shaibani, Shanya Sharma, Urmish Thakker, Khalid Almubarak, Xiangru Tang, Dragomir Radev, Mike Tian-jian Jiang, and Alexander Rush. 2022. PromptSource: An integrated development environment and repository for natural language prompts. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics: System Demonstrations, pages 93–104, Dublin, Ireland. Association for Computational Linguistics.
  • Bandarkar et al. (2024) Lucas Bandarkar, Davis Liang, Benjamin Muller, Mikel Artetxe, Satya Narayan Shukla, Donald Husa, Naman Goyal, Abhinandan Krishnan, Luke Zettlemoyer, and Madian Khabsa. 2024. The belebele benchmark: a parallel reading comprehension dataset in 122 language variants. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 749–775, Bangkok, Thailand. Association for Computational Linguistics.
  • Barbaresi (2021) Adrien Barbaresi. 2021. Trafilatura: A web scraping library and command-line tool for text discovery and extraction. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing: System Demonstrations, pages 122–131, Online. Association for Computational Linguistics.
  • Baucells et al. (2025) Irene Baucells, Javier Aula-Blasco, Iria de Dios-Flores, Silvia Paniagua Suárez, Naiara Perez, Anna Salles, Susana Sotelo Docio, Júlia Falcão, Jose Javier Saiz, Robiert Sepulveda Torres, Jeremy Barnes, Pablo Gamallo, Aitor Gonzalez-Agirre, German Rigau, and Marta Villegas. 2025. IberoBench: A benchmark for LLM evaluation in Iberian languages. In Proceedings of the 31st International Conference on Computational Linguistics, pages 10491–10519, Abu Dhabi, UAE. Association for Computational Linguistics.
  • Burchell et al. (2023) Laurie Burchell, Alexandra Birch, Nikolay Bogoychev, and Kenneth Heafield. 2023. An open dataset and model for language identification. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 865–879, Toronto, Canada. Association for Computational Linguistics.
  • Burchell et al. (2025) Laurie Burchell, Ona De Gibert Bonet, Nikolay Arefyev, Mikko Aulamo, Marta Bañón, Pinzhen Chen, Mariia Fedorova, Liane Guillou, Barry Haddow, Jan Hajič, Jindřich Helcl, Erik Henriksson, Mateusz Klimaszewski, Ville Komulainen, Andrey Kutuzov, Joona Kytöniemi, Veronika Laippala, Petter Mæhlum, Bhavitvya Malik, Farrokh Mehryary, Vladislav Mikhailov, Nikita Moghe, Amanda Myntti, Dayyán O’Brien, Stephan Oepen, Proyag Pal, Jousia Piha, Sampo Pyysalo, Gema Ramírez-Sánchez, David Samuel, Pavel Stepachev, Jörg Tiedemann, Dušan Variš, Tereza Vojtěchová, and Jaume Zaragoza-Bernabeu. 2025. An expanded massive multilingual dataset for high-performance language technologies (HPLT). In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 17452–17485, Vienna, Austria. Association for Computational Linguistics.
  • Burchell et al. (2024) Laurie Burchell, Jean Maillard, Antonios Anastasopoulos, Christian Federmann, Philipp Koehn, and Skyler Wang. 2024. Findings of the WMT 2024 shared task of the open language data initiative. In Proceedings of the Ninth Conference on Machine Translation, pages 110–117, Miami, Florida, USA. Association for Computational Linguistics.
  • de Gibert et al. (2025) Ona de Gibert, Joseph Attieh, Teemu Vahtola, Mikko Aulamo, Zihao Li, Raúl Vázquez, Tiancheng Hu, and Jörg Tiedemann. 2025. Scaling low-resource mt via synthetic data generation with llms. arXiv preprint arXiv:2505.14423.
  • de Gibert et al. (2024) Ona de Gibert, Graeme Nail, Nikolay Arefyev, Marta Bañón, Jelmer van der Linde, Shaoxiong Ji, Jaume Zaragoza-Bernabeu, Mikko Aulamo, Gema Ramírez-Sánchez, Andrey Kutuzov, Sampo Pyysalo, Stephan Oepen, and Jörg Tiedemann. 2024. A new massive multilingual dataset for high-performance language technologies. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 1116–1128, Torino, Italia. ELRA and ICCL.
  • Dementieva et al. (2024) Daryna Dementieva, Valeriia Khylenko, Nikolay Babakov, and Georg Groh. 2024. Toxicity classification in Ukrainian. In Proceedings of the 8th Workshop on Online Abuse and Harms (WOAH 2024), pages 244–255, Mexico City, Mexico. Association for Computational Linguistics.
  • Deutsch et al. (2025a) Daniel Deutsch, Eleftheria Briakou, Isaac Caswell, Mara Finkelstein, Rebecca Galor, Juraj Juraska, Geza Kovacs, Alison Lui, Ricardo Rei, Jason Riesa, et al. 2025a. Wmt24++: Expanding the language coverage of wmt24 to 55 languages & dialects. arXiv preprint arXiv:2502.12404.
  • Deutsch et al. (2025b) Daniel Deutsch, Eleftheria Briakou, Isaac Rayburn Caswell, Mara Finkelstein, Rebecca Galor, Juraj Juraska, Geza Kovacs, Alison Lui, Ricardo Rei, Jason Riesa, Shruti Rijhwani, Parker Riley, Elizabeth Salesky, Firas Trabelsi, Stephanie Winkler, Biao Zhang, and Markus Freitag. 2025b. WMT24++: Expanding the language coverage of WMT24 to 55 languages & dialects. In Findings of the Association for Computational Linguistics: ACL 2025, pages 12257–12284, Vienna, Austria. Association for Computational Linguistics.
  • Doshi et al. (2024) Meet Doshi, Raj Dabre, and Pushpak Bhattacharyya. 2024. Pretraining language models using translationese. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages 5843–5862, Miami, Florida, USA. Association for Computational Linguistics.
  • Fajcik et al. (2025) Martin Fajcik, Martin Docekal, Jan Dolezal, Karel Ondrej, Karel Beneš, Jan Kapsa, Pavel Smrz, Alexander Polok, Michal Hradis, Zuzana Neverilova, et al. 2025. BenCzechMark: A Czech-centric Multitask and Multimetric Benchmark for Large Language Models with Duel Scoring Mechanism. Transactions of the Association for Computational Linguistics, 13:1068–1095.
  • Faysse et al. (2024) Manuel Faysse, Patrick Fernandes, Nuno M Guerreiro, António Loison, Duarte Miguel Alves, Caio Corro, Nicolas Boizard, João Alves, Ricardo Rei, Pedro Henrique Martins, et al. 2024. CroissantLLM: A Truly Bilingual French-English Language Model. Transactions on Machine Learning Research.
  • Fernandes et al. (2023) Patrick Fernandes, Kayo Yin, Emmy Liu, André Martins, and Graham Neubig. 2023. When does translation require context? a data-driven, multilingual exploration. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 606–626, Toronto, Canada. Association for Computational Linguistics.
  • Fourrier et al. (2024) Clémentine Fourrier, Nathan Habib, Alina Lozovskaya, Konrad Szafer, and Thomas Wolf. 2024. Open llm leaderboard v2.
  • Gao et al. (2024) Leo Gao, Jonathan Tow, Baber Abbasi, Stella Biderman, Sid Black, Anthony DiPofi, Charles Foster, Laurence Golding, Jeffrey Hsu, Alain Le Noac’h, Haonan Li, Kyle McDonell, Niklas Muennighoff, Chris Ociepa, Jason Phang, Laria Reynolds, Hailey Schoelkopf, Aviya Skowron, Lintang Sutawika, Eric Tang, Anish Thite, Ben Wang, Kevin Wang, and Andy Zou. 2024. The language model evaluation harness.
  • Ivanyuk-Skulskiy et al. (2021) Bogdan Ivanyuk-Skulskiy, Anton Zaliznyi, Oleksand Reshetar, Oleksiy Protsyk, Bohdan Romanchuk, and Vladyslav Shpihanovych. 2021. ua_datasets: a collection of ukrainian language datasets.
  • Jentoft and Samuel (2023) Matias Jentoft and David Samuel. 2023. NoCoLA: The Norwegian corpus of linguistic acceptability. In Proceedings of the 24th Nordic Conference on Computational Linguistics (NoDaLiDa), pages 610–617, Tórshavn, Faroe Islands. University of Tartu Library.
  • Jumelet et al. (2025) Jaap Jumelet, Leonie Weissweiler, Joakim Nivre, and Arianna Bisazza. 2025. Multiblimp 1.0: A massively multilingual benchmark of linguistic minimal pairs.
  • Junczys-Dowmunt et al. (2018) Marcin Junczys-Dowmunt, Roman Grundkiewicz, Tomasz Dwojak, Hieu Hoang, Kenneth Heafield, Tom Neckermann, Frank Seide, Ulrich Germann, Alham Fikri Aji, Nikolay Bogoychev, André F. T. Martins, and Alexandra Birch. 2018. Marian: Fast neural machine translation in C++. In Proceedings of ACL 2018, System Demonstrations, pages 116–121, Melbourne, Australia. Association for Computational Linguistics.
  • Kudugunta et al. (2023) Sneha Kudugunta, Isaac Caswell, Biao Zhang, Xavier Garcia, Derrick Xin, Aditya Kusupati, Romi Stella, Ankur Bapna, and Orhan Firat. 2023. MADLAD-400: A multilingual and document-level large audited dataset. In Proceedings of the 37th International Conference on Neural Information Processing Systems, NeurIPS ’23, Red Hook, NY, USA. Curran Associates Inc.
  • Lee et al. (2022) Katherine Lee, Daphne Ippolito, Andrew Nystrom, Chiyuan Zhang, Douglas Eck, Chris Callison-Burch, and Nicholas Carlini. 2022. Deduplicating training data makes language models better. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 8424–8445, Dublin, Ireland. Association for Computational Linguistics.
  • Lozhkov et al. (2024) Anton Lozhkov, Loubna Ben Allal, Leandro von Werra, and Thomas Wolf. 2024. Fineweb-edu: the finest collection of educational content.
  • Luukkonen et al. (2023) Risto Luukkonen, Ville Komulainen, Jouni Luoma, Anni Eskelinen, Jenna Kanerva, Hanna-Mari Kupari, Filip Ginter, Veronika Laippala, Niklas Muennighoff, Aleksandra Piktus, Thomas Wang, Nouamane Tazi, Teven Scao, Thomas Wolf, Osma Suominen, Samuli Sairanen, Mikko Merioksa, Jyrki Heinonen, Aija Vahtola, Samuel Antao, and Sampo Pyysalo. 2023. FinGPT: Large generative models for a small language. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 2710–2726, Singapore. Association for Computational Linguistics.
  • Merx et al. (2025) Raphaël Merx, Hanna Suominen, Trevor Cohn, and Ekaterina Vylomova. 2025. Openwho: A document-level parallel corpus for health translation in low-resource languages. arXiv preprint arXiv:2508.16048.
  • Mikhailov et al. (2025) Vladislav Mikhailov, Tita Enstad, David Samuel, Hans Christian Farsethås, Andrey Kutuzov, Erik Velldal, and Lilja Øvrelid. 2025. NorEval: A Norwegian language understanding and generation evaluation benchmark. In Findings of the Association for Computational Linguistics: ACL 2025, pages 3495–3541, Vienna, Austria. Association for Computational Linguistics.
  • Muennighoff et al. (2023a) Niklas Muennighoff, Alexander M Rush, Boaz Barak, Teven Le Scao, Nouamane Tazi, Aleksandra Piktus, Sampo Pyysalo, Thomas Wolf, and Colin Raffel. 2023a. Scaling Data-Constrained Language Models. In Thirty-seventh Conference on Neural Information Processing Systems.
  • Muennighoff et al. (2023b) Niklas Muennighoff, Thomas Wang, Lintang Sutawika, Adam Roberts, Stella Biderman, Teven Le Scao, M Saiful Bari, Sheng Shen, Zheng Xin Yong, Hailey Schoelkopf, Xiangru Tang, Dragomir Radev, Alham Fikri Aji, Khalid Almubarak, Samuel Albanie, Zaid Alyafeai, Albert Webson, Edward Raff, and Colin Raffel. 2023b. Crosslingual generalization through multitask finetuning. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 15991–16111, Toronto, Canada. Association for Computational Linguistics.
  • Myntti et al. (2024) Amanda Myntti, Liina Repo, Elian Freyermuth, Antti Kanner, Veronika Laippala, and Erik Henriksson. 2024. Intersecting register and genre: Understanding the contents of web-crawled corpora. In Proceedings of the 4th International Conference on Natural Language Processing for Digital Humanities, pages 386–397, Miami, USA. Association for Computational Linguistics.
  • O’Brien et al. (2025) Dayyán O’Brien, Bhavitvya Malik, Ona de Gibert, Pinzhen Chen, Barry Haddow, and Jörg Tiedemann. 2025. Dochplt: A massively multilingual document-level translation dataset. arXiv preprint arXiv:2508.13079.
  • Penedo et al. (2024) Guilherme Penedo, Hynek Kydlíček, Anton Lozhkov, Margaret Mitchell, Colin A Raffel, Leandro Von Werra, Thomas Wolf, et al. 2024. The FineWeb datasets: Decanting the web for the finest text data at scale. Advances in Neural Information Processing Systems, 37:30811–30849.
  • Penedo et al. (2025) Guilherme Penedo, Hynek Kydlíček, Vinko Sabolčec, Bettina Messmer, Negar Foroutan, Amir Hossein Kargaran, Colin Raffel, Martin Jaggi, Leandro Von Werra, and Thomas Wolf. 2025. FineWeb2: One pipeline to scale them all – adapting pre-training data processing to every language.
  • Pezeshkpour and Hruschka (2024) Pouya Pezeshkpour and Estevam Hruschka. 2024. Large language models sensitivity to the order of options in multiple-choice questions. In Findings of the Association for Computational Linguistics: NAACL 2024, pages 2006–2017, Mexico City, Mexico. Association for Computational Linguistics.
  • Qorib et al. (2025) Muhammad Reza Qorib, Junyi Li, and Hwee Tou Ng. 2025. Just go parallel: Improving the multilingual capabilities of large language models. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 33411–33424, Vienna, Austria. Association for Computational Linguistics.
  • Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. 2020. Exploring the limits of transfer learning with a unified text-to-text transformer. Journal of Machine Learning Research, 21(140):1–67.
  • Rahimi et al. (2019) Afshin Rahimi, Yuan Li, and Trevor Cohn. 2019. Massively multilingual transfer for NER. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 151–164, Florence, Italy. Association for Computational Linguistics.
  • Ramos et al. (2025) Miguel Moura Ramos, Patrick Fernandes, Sweta Agrawal, and André FT Martins. 2025. Multilingual contextualization of large language models for document-level machine translation. arXiv preprint arXiv:2504.12140.
  • Rofin et al. (2023) Mark Rofin, Vladislav Mikhailov, Mikhail Florinsky, Andrey Kravchenko, Tatiana Shavrina, Elena Tutubalina, Daniel Karabekyan, and Ekaterina Artemova. 2023. Vote’n’rank: Revision of benchmarking with social choice theory. In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics, pages 670–686, Dubrovnik, Croatia. Association for Computational Linguistics.
  • Romanou et al. (2025) Angelika Romanou, Negar Foroutan, Anna Sotnikova, Sree Harsha Nelaturu, Shivalika Singh, Rishabh Maheshwary, Micol Altomare, Zeming Chen, Mohamed A. Haggag, Snegha A, Alfonso Amayuelas, Azril Hafizi Amirudin, Danylo Boiko, Michael Chang, Jenny Chim, Gal Cohen, Aditya Kumar Dalmia, Abraham Diress, Sharad Duwal, Daniil Dzenhaliou, Daniel Fernando Erazo Florez, Fabian Farestam, Joseph Marvin Imperial, Shayekh Bin Islam, Perttu Isotalo, Maral Jabbarishiviari, Börje F. Karlsson, Eldar Khalilov, Christopher Klamm, Fajri Koto, Dominik Krzemiński, Gabriel Adriano de Melo, Syrielle Montariol, Yiyang Nan, Joel Niklaus, Jekaterina Novikova, Johan Samir Obando Ceron, Debjit Paul, Esther Ploeger, Jebish Purbey, Swati Rajwal, Selvan Sunitha Ravi, Sara Rydell, Roshan Santhosh, Drishti Sharma, Marjana Prifti Skenduli, Arshia Soltani Moakhar, Bardia soltani moakhar, Ayush Kumar Tarun, Azmine Toushik Wasi, Thenuka Ovin Weerasinghe, Serhan Yilmaz, Mike Zhang, Imanol Schlag, Marzieh Fadaee, Sara Hooker, and Antoine Bosselut. 2025. INCLUDE: Evaluating multilingual language understanding with regional knowledge. In The Thirteenth International Conference on Learning Representations.
  • Romanyshyn et al. (2024) Mariana Romanyshyn, Oleksiy Syvokon, and Roman Kyslyi. 2024. The UNLP 2024 shared task on fine-tuning large language models for Ukrainian. In Proceedings of the Third Ukrainian Natural Language Processing Workshop (UNLP) @ LREC-COLING 2024, pages 67–74, Torino, Italia. ELRA and ICCL.
  • Samuel et al. (2023) David Samuel, Andrey Kutuzov, Samia Touileb, Erik Velldal, Lilja Øvrelid, Egil Rønningstad, Elina Sigdel, and Anna Palatkina. 2023. NorBench – a benchmark for Norwegian language models. In Proceedings of the 24th Nordic Conference on Computational Linguistics (NoDaLiDa), pages 618–633, Tórshavn, Faroe Islands. University of Tartu Library.
  • Sclar et al. (2024) Melanie Sclar, Yejin Choi, Yulia Tsvetkov, and Alane Suhr. 2024. Quantifying Language Models’ Sensitivity to Spurious Features in Prompt Design or: How I learned to start worrying about prompt formatting. In The Twelfth International Conference on Learning Representations.
  • Shoeybi et al. (2019) Mohammad Shoeybi, Mostofa Patwary, Raul Puri, Patrick LeGresley, Jared Casper, and Bryan Catanzaro. 2019. Megatron-LM: Training Multi-Billion Parameter Language Models Using Model Parallelism. arXiv preprint arXiv:1909.08053.
  • Singh et al. (2025) Shivalika Singh, Angelika Romanou, Clémentine Fourrier, David Ifeoluwa Adelani, Jian Gang Ngui, Daniel Vila-Suero, Peerat Limkonchotiwat, Kelly Marchisio, Wei Qi Leong, Yosephine Susanto, Raymond Ng, Shayne Longpre, Sebastian Ruder, Wei-Yin Ko, Antoine Bosselut, Alice Oh, Andre Martins, Leshem Choshen, Daphne Ippolito, Enzo Ferrante, Marzieh Fadaee, Beyza Ermis, and Sara Hooker. 2025. Global MMLU: Understanding and addressing cultural and linguistic biases in multilingual evaluation. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 18761–18799, Vienna, Austria. Association for Computational Linguistics.
  • Su et al. (2025) Dan Su, Kezhi Kong, Ying Lin, Joseph Jennings, Brandon Norick, Markus Kliegl, Mostofa Patwary, Mohammad Shoeybi, and Bryan Catanzaro. 2025. Nemotron-CC: Transforming Common Crawl into a refined long-horizon pretraining dataset. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 2459–2475, Vienna, Austria. Association for Computational Linguistics.
  • Tay et al. (2023) Yi Tay, Mostafa Dehghani, Vinh Q. Tran, Xavier Garcia, Jason Wei, Xuezhi Wang, Hyung Won Chung, Dara Bahri, Tal Schuster, Steven Zheng, Denny Zhou, Neil Houlsby, and Donald Metzler. 2023. UL2: Unifying language learning paradigms. In The Eleventh International Conference on Learning Representations.
  • Team (2025) Gemma Team. 2025. Gemma 3. Technical report, Google.
  • Team et al. (2022) NLLB Team, Marta R. Costa-jussà, James Cross, Onur Çelebi, Maha Elbayad, Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht, Jean Maillard, Anna Sun, Skyler Wang, Guillaume Wenzek, Al Youngblood, Bapi Akula, Loic Barrault, Gabriel Mejia Gonzalez, Prangthip Hansanti, John Hoffman, Semarley Jarrett, Kaushik Ram Sadagopan, Dirk Rowe, Shannon Spruit, Chau Tran, Pierre Andrews, Necip Fazil Ayan, Shruti Bhosale, Sergey Edunov, Angela Fan, Cynthia Gao, Vedanuj Goswami, Francisco Guzmán, Philipp Koehn, Alexandre Mourachko, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, and Jeff Wang. 2022. No language left behind: Scaling human-centered machine translation.
  • Tiedemann et al. (2023) Jörg Tiedemann, Mikko Aulamo, Daria Bakshandaeva, Michele Boggia, Stig-Arne Grönroos, Tommi Nieminen, Alessandro Raganato, Yves Scherrer, Raul Vazquez, and Sami Virpioja. 2023. Democratizing neural machine translation with OPUS-MT. Language Resources and Evaluation, pages 1–43.
  • Tiedemann and de Gibert (2023) Jörg Tiedemann and Ona de Gibert. 2023. The OPUS-MT dashboard – a toolkit for a systematic evaluation of open machine translation models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations), pages 315–327, Toronto, Canada. Association for Computational Linguistics.
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971.
  • Üstün et al. (2024) Ahmet Üstün, Viraat Aryabumi, Zheng Yong, Wei-Yin Ko, Daniel D’souza, Gbemileke Onilude, Neel Bhandari, Shivalika Singh, Hui-Lee Ooi, Amr Kayid, Freddie Vargus, Phil Blunsom, Shayne Longpre, Niklas Muennighoff, Marzieh Fadaee, Julia Kreutzer, and Sara Hooker. 2024. Aya model: An instruction finetuned open-access multilingual language model. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 15894–15939, Bangkok, Thailand. Association for Computational Linguistics.
  • Wang et al. (2025) Jiayi Wang, Yao Lu, Maurice Weber, Max Ryabinin, David Adelani, Yihong Chen, Raphael Tang, and Pontus Stenetorp. 2025. Multilingual language model pretraining using machine-translated data.
  • Xue et al. (2021) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2021. mT5: A massively multilingual pre-trained text-to-text transformer. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 483–498, Online. Association for Computational Linguistics.
  • Zhang et al. (2025) Biao Zhang, Fedor Moiseev, Joshua Ainslie, Paul Suganthan, Min Ma, Surya Bhupatiraju, Fede Lebron, Orhan Firat, Armand Joulin, and Zhe Dong. 2025. Encoder-decoder gemma: Improving the quality-efficiency trade-off via adaptation.
  • Zhang et al. (2024) Yuanchi Zhang, Yile Wang, Zijun Liu, Shuo Wang, Xiaolong Wang, Peng Li, Maosong Sun, and Yang Liu. 2024. Enhancing multilingual capabilities of large language models through self-distillation from resource-rich languages. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 11189–11204.